a double slit interference pattern is shown below

A double slit interference pattern is shown below.
Answer #1

for double slit inteference

the position of minima is given by y = m*lambda*R/2d


m = 1 , 3, 5 ....

for first order minima m = 1


given y = 4.8*10^-3 m

R = distance of screen from slit = 2 m


lambda = wavelength = 633*10^-9 m

d = slits seperation


4.8*10^-3 = 1*633*10^-9*2/(2*d)

d = 1.318*10^-4 m

===================

for single slit diffraction

y = m*lambda/a

here m = 6 the minima


a = slit width


4.8*10^-3 = 6*633*10^-9/a


a = 7.9125*10^-4 m